OpenStudy (anonymous):

simplify the radical expression square root of 24a^4b^5c^9 show work

OpenStudy (anonymous):

think of them individually such as sqrt (24a^4) and sqrt (b^5) and sqrt (c^9)

OpenStudy (anonymous):

i need help with radicals period

OpenStudy (anonymous):

sqrt (24a^4) =2a^2 x sqrt (6)

OpenStudy (amistre64):

turn it into an exponent; simply the exponent and redress it as a radical perhaps?

OpenStudy (anonymous):

sorry about my post - it belonged to another question and i cant delete it

OpenStudy (anonymous):

its okay i just need a lot of help

OpenStudy (amistre64):

\[\sqrt{24a^4b^5c^9 }\] \[24^{1/2}\ a^{4/2}\ b^{5/2}\ c^{9/2} \] \[4^{1/2}6^{1/2}\ a^{2}\ b^{4/2+1/2}\ c^{8/2 + 1/2} \] \[2\sqrt{6}\ a^{2}\ b^{2}\ b^{1/2}\ c^{4}c^{1/2} \] \[2\sqrt{6}\ a^{2}\ b^{2}\ \sqrt{b}\ c^{4}\sqrt{c} \] \[2a^2b^2c^4\sqrt{6bc} \] maybe

OpenStudy (amistre64):

most of that nonesense is just for practice :)

OpenStudy (anonymous):

i got \[2a^2b^2c^4\sqrt{6bc}\] is this correct?